OpenStudy (anonymous):

solve for cot(2theta/3)=squareroot of 3

OpenStudy (anonymous):

If you realize mebbe that cot(n)=1/tan(n), and thus tan(2θ/3)=1/√(3), 2θ/3=arctan(1/√(3)), or θ=(3/2)arctan(1/√(3)). Which should be solvable via the miracle known as the calculator.

OpenStudy (anonymous):

we r not allowed to use calculator

OpenStudy (anonymous):

Awright. The angle is 30 degrees. Memorize that.

OpenStudy (anonymous):

what can u show me step by step please

OpenStudy (anonymous):

cot(2θ/3)=√(3); solve for θ 1/tan(2θ/3)=√(3); definition of cotangent tan(2θ/3)=1/√(3); raise both sides to the power of -1 2θ/3=arctan(1/√(3)); take the arctan of both sides θ=(3/2)arctan(1/√(3)); multiply both sides by 3/2 θ=(3/2)(π/6); remember that arctan(1/√(3)) is (π/6) or 30° θ=π/4

OpenStudy (anonymous):

The last step is obviously a simplification.

OpenStudy (anonymous):

what is arctan

OpenStudy (anonymous):

The inverse tangent; it's easier to write it as such on OpenStudy.

OpenStudy (anonymous):

so arctan is the same as cotangent?

OpenStudy (anonymous):

No. Let's say I had y=tan(x), okay? 1/y=1/tan(x)=cot(x). However, x=arctan(y).

OpenStudy (anonymous):

is there another way to solve without using arctan cause im supposed to use the trig identities

OpenStudy (anonymous):

Yes, there's another way without trig identities. It looks exactly the same, except instead of taking the arctan of both sides, you pretend there's no arctan, and just memorize all the steps in between.

OpenStudy (anonymous):

no with identities and without arctan

OpenStudy (anonymous):

Arctan is a basic trig identity. It's the functional inverse of tan. As I said, if your teacher hasn't mentioned arctan yet, it's because s/he wants you to do it WITH arctan but pretend it isn't there.

OpenStudy (anonymous):

okay